OpenStudy (aaronandyson):

Solve using formula 5(x^2 - 1) = 3x.Give your answer correct to 2 significant figures. @iGreen @confluxepic @Conqueror @AnswerMyQuestions

OpenStudy (aaronandyson):

@chosenmatt

OpenStudy (aaronandyson):

@sleepyjess

OpenStudy (aaronandyson):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[5(x^2 - 1) = 3x\]\]

OpenStudy (aaronandyson):

yes

OpenStudy (misty1212):

start with \[5x^2-5=3x\]

OpenStudy (misty1212):

then set it equal to zero and get \[5x^2-3x-5=0\]

OpenStudy (aaronandyson):

now?

OpenStudy (misty1212):

use the quadratic formula i guess, since this one does not factor

OpenStudy (misty1212):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \[a=5,b=-3,c=-5\]

OpenStudy (misty1212):

if that is too much work, you can always cheat you are going to need a calculator anyway, since they are asking for the answer as a decimal http://www.wolframalpha.com/input/?i=5x%5E2-3x-5%3D0

OpenStudy (aaronandyson):

We are not allowed to use a calculator.

OpenStudy (misty1212):

scroll down to where you see "solutions" and click on "approximate form"

OpenStudy (misty1212):

then you cannot answer it it asks "Give your answer correct to 2 significant figures. "

OpenStudy (misty1212):

that means you HAVE TO use a calculator otherwise you cannot compute the radical

OpenStudy (misty1212):

the answer without a calculator should be \[\frac{3\pm\sqrt{109}}{10}\] but if you want a decimal you need a calculator